OpenStudy (anonymous):

Find the area of polygon. Round to the nearest tenth. a regular octagon with a perimeter of 72 inches

OpenStudy (anonymous):

@shreya.

OpenStudy (anonymous):

area of an octagon is \[2\cot \frac{ \pi }{ 8 }a^{2}\]

OpenStudy (anonymous):

the diameter wold be same as the length side

OpenStudy (anonymous):

where a is the side of the octagon which in this case it's 9

OpenStudy (anonymous):

what do u mean ??

OpenStudy (anonymous):

i must use this formula A=1/2(a)(P)

OpenStudy (anonymous):

yes,I mean the radius wold be 9 or no?

OpenStudy (anonymous):

you mean the circumradius ?

OpenStudy (anonymous):



OpenStudy (anonymous):

no 9 is the side of the octagon

OpenStudy (anonymous):

 i need to find the apothem